# Dynamics in an emergent quantum-like state space generated by a nonlinear classical network

### v1

@@ V.2 Synchronization
We now describe numerical results for a network of phase oscillators that are connected in a way prescribed by the product of two QL bits. For these calculations, we have performed an average over initial phases as well as an average over 500 graphs (see Methods). First, we set the initial phase distribution to be, essentially, the uniform distribution and the coupling to be relatively strong, $K=250$. These conditions allow us to observe synchronization of the oscillators. The line in Fig. 3a shows the ensemble average order parameter, Eq.11, predicted by propagating Eqs. 10 over 80 periods of the mean natural frequency of the oscillators. At early time the order parameter is small because the initial condition specifies a near uniform phase distribution. As time increases, the order parameter increases to approximately 1, showing that the nonlinear evolution of the phase oscillator system synchronizes the oscillator phases.
@@
Concomitantly, the state associated with the greatest eigenvalue in the spectrum of the state space represented by the graph becomes less mixed. The state evolves in the effective state space by Eq. 26, according to the time dependence of the coefficients. We highlight this result for the time points labeled numerically in Fig. 3a. Corresponding spectra are all identical, as expected. An example is plotted in Fig. 3b. Purity of the state with greatest eigenvalue is indicated together with the time point, written as a fraction of the total time of the simulation. For example, $\tfrac{1}{8}T$ equates to 10 periods of oscillation. Note that the purity for a fully mixed system of dimension 4, with a diagonal density matrix, is 0.25 (see Sec. 4d of Ref. [58]). The initial purity of 0.25 means that the state is mixed because we have no information about the phases of the oscillators. As the phases synchronize, the variance of the phases diminishes, and the purity increases. The relationship between phase distribution and purity is discussed in Ref. [53].
@@
We show this purity-preserving regime in Fig. 4. After an initial period when the oscillators synchronize, the system settles to equilibrium dynamics where the oscillators are locked in a limit cycle, that is, a closed, isolated (stable) phase trajectory. Owing to the way we define the oscillator network of each QL bit, these stable dynamics involve classical superpositions of oscillators local to each subgraph. In the corresponding state space, we find that when the classical system is synchronized, the purity of the emergent state is approximately 1. The observed dynamics in the state space are purity-preserving. In this regime, dynamics that ensue from abruptly changing basis, applying a driving field, imposing quantum quench conditions, and so on, can be unitary. To explore this regime in future work, it would be useful define the properties of the QL state graphs a little differently. For example, we can assign a frequency $\omega_{1}$ to vertices in the subgraph $G_{a1}$ and $\omega_{2}\neq\omega_{1}$ to vertices in the subgraph $G_{a2}$, thereby allowing the two states of the QL bit to have a frequency gap.
@@
Figure 4: (a) Ensemble-averaged Kuramoto order parameter as a function of time predicted for the system of oscillators. The system equilibrates to limit cycle dynamics, indicating a synchronized classical system. (b) The state space concomitantly evolves to an equilibrium of purity-preserving dynamics.
@@
In the section, we have generalized our prior model, allowing any phase for the edges. We generalized the model further by explicitly considering the degree of phase synchronization in the classical network. The extent of synchronization of the underlying classical network controls the mixedness of the emergent states. Hence, this property of the variables defined by the classical network influences an important property of the states—phase coherence of the basis states is essential for sustaining stable superposition states. Depending on the properties of the of the classical phase oscillator network, the QL states can exhibit purity-preserving dynamics, or they can be completely mixed and undergo classical dynamics, or any case in between. When the oscillator network is synchronized, that is, the phases are coherent, then the QL states have high purity. Physically, this phase coherence is similar to properties of a wave-like system underlying the states.
V.3 Decoherence
We have focused on synchronization of the classical network to produce robust QL states. Conversely, we hypothesize that a poorly synchronized network will decohere a pure QL state associated with an initially synchronized classical network. That is, the opposite dynamics to those illustrated in Fig. 4 will ensue. An obvious example is a network of uncoupled oscillators with a frequency distribution. Starting in-phase, free evolution of the oscillators will dephase the order parameter and states in the network.
@@ V.3 Decoherence
Figure 5: Results of calculations of an oscillator system comprising the Cartesian product of two QL bits, using the same procedures and parameters as the previous calculations, except here we set the coupling to be weak, $K=30$ and the initial condition sets all the oscillators in phase. Here the averaging is carried out over a single graph. (a) The order parameter, initially 1, indicating the initial synchronization condition, decays as the oscillators become out of phase. (b) As the oscillators dephase, the purity of the emergent state decays to an equilibrium mixed state. (c) Concomitantly, off-diagonal elements of the density matrix in the product basis decay to zero, highlighting loss of superpositions.
@@
In Fig. 5 we show results of calculations of an oscillator system comprising the Cartesian product of two QL bits, using most of the same procedures and parameters as the previous calculations. Two key parameters were changed: the coupling was set to be weak, $K=30$, and the initial condition sets all the oscillators in phase. We show results from two similar sets of calculations. One where we average only over the initial phase conditions. In the other calculation we average also over the ensemble of graphs. In each case, the average is over 500 realizations of the system.
@@
V.4 Discussion
@@
The model studied here suggests a way to understand a smooth QL to classical transition in the state space. This would be observed if the propensity of the classical network to synchronize diminishes as the size of the system increases. An open question, therefore, is whether (and how) there can be a distance dependence governing the decoherence so that distant superpositions naturally collapse. For example, we can imagine that the noise that gives the oscillators the frequency distribution is correlated on molecular length scales, but uncorrelated on longer length scales. This setup might serve to collapse distant superpositions, but preserve superpositions among closely-spaced systems.
@@
Thinking along these lines could enable a connection to be made to work aiming for a unified theory for the dynamics of microscopic and macroscopic systems, in particular, aiming to explain a mechanism for how superposition states can spontaneously collapse[3, 35, 34, 65]. Those studies incorporate a term into the Schrödinger equation that has no effect for typical quantum states, but for distant superpositions a stochastic, nonlinear term emerges that can initiate collapse of superpositions. The model considered in the present work appears to produce a similar result. We posited a somewhat different mechanism than proposed in prior important papers[3, 35, 34, 65] because in the present work the nonlinearity comes from the dynamics of the explicit classical system underlying the state space and it only controls phases of the underlying variables. The crucial parameter that impacts the states is the phase coherence of the classical network. However, the net result would be a modification of the Schrödinger equation prediction of dynamics in the state space. Indeed, we have noted how, when the oscillator network is poorly synchronized, it acts as a kind of bath for the state space. Thereby, it introduces a mechanism for decoherence similar to models for open quantum systems[66], where a system interacts with an external bath comprising many degrees of freedom. There is therefore an interesting connection between the models.
@@ VI Conclusions
We studied a specific example of a large classical system that maps, via a graph, to a state space that we call quantum like (QL) because it mimics the kinds of states typical of quantum systems. The classical system is a network of coupled phase oscillators, of the kind extensively studied in a wide variety of contexts. Such networks of coupled phase oscillator evolve according to the nonlinear dynamics associated with synchronization of the oscillator phases. We reported numerical simulations for dynamics of this system, starting from an initial condition of a random distribution of oscillator phases. The network topology was blueprinted by a graph that generates the QL states. We used the phases of the classical system to weight the edges of the QL state graph at various time points during the dynamics, then mapped those graphs to the QL state space. Studying the state space, we constructed the density matrix in an effective basis of two coupled QL bits and showed that the mixedness of the emergent state diminished as the underlying classical system synchronized.
@@
Key conclusions of the work are that nonlinear dynamics in the classical system controls the dynamics in the state space. The dynamics can help quantum-like states be robust to noise and disorder, or they can act like an intrinsic environment to produce decoherence in the state space. We showed numerical studies of both scenarios. Nevertheless, these dynamics convey linear maps in the state space (Prop. 1). This setup suggests a mechanism by which underlying variables act to collapse superpositions between product basis states in the QL system. The concept suggests a plausible mechanism that could potentially be generalized to explain the quantum-to-classical transition.

### v3

@@ V.4 Synchronization
We now describe numerical results for a network of phase oscillators that are connected in a way prescribed by the product of two QL bits. For these calculations, we have performed an average over initial phases (see Sec. V.B). First, we set the initial phase distribution to be, essentially, the uniform distribution and the coupling to be relatively strong, $K=250$. These conditions allow us to observe synchronization of the oscillators. The line in Fig. 3a shows the ensemble average order parameter, Eq.11, predicted by propagating Eqs. 10 over 80 periods of the mean natural frequency of the oscillators. At early time the order parameter is small because the initial condition specifies a near uniform phase distribution. As time increases, the order parameter increases to approximately 1, showing that the nonlinear evolution of the phase oscillator system synchronizes the oscillator phases.
@@
Concomitantly, the state associated with the greatest eigenvalue in the spectrum of the state space represented by the graph becomes less mixed. The state evolves in the effective state space according to the time dependence of the coefficients in Eq. 25. We highlight this result for the time points labeled numerically in Fig. 3a. Corresponding spectra are all identical, as expected. An example is plotted in Fig. 3b. The purity of the state with greatest eigenvalue is indicated together with the time point, written as a fraction of the total time of the simulation. For example, $\tfrac{1}{8}T$ equates to 10 periods of oscillation. Note that the purity for a fully mixed system of dimension 4, with a diagonal density matrix, is 0.25 (see Sec. 4d of Ref. [49]). The initial purity of 0.25 means that the state is mixed because we have no information about the phases of the oscillators. As the phases synchronize, the variance of the phases diminishes, and the purity increases. The relationship between phase distribution and purity is discussed in Ref. [40].
@@
We show this purity-preserving regime in Fig. 4. In the figure, calculations with various parameters are compared, as specified in the caption. Each simulation was averaged over at least 100 state preparations. After an initial period when the oscillators synchronize, the system settles to equilibrium dynamics. When $K$ is sufficiently large the oscillators are locked in a limit cycle, that is, a closed, isolated (stable) phase trajectory. Owing to the way we define the oscillator network of each QL bit, these stable dynamics involve classical superpositions of oscillators local to each subgraph. In the corresponding state space, we find that when the classical system is synchronized, the purity of the emergent QL state is approximately 1. The observed dynamics in the state space are purity-preserving. In this regime, dynamics that ensue from abruptly changing basis, applying a driving field, imposing quantum quench conditions, and so on, can be unitary. To explore this regime in future work, it would be useful define the properties of the QL state graphs a little differently. For example, we can assign a frequency $\omega_{1}$ to vertices in the subgraph $G_{a1}$ and $\omega_{2}\neq\omega_{1}$ to vertices in the subgraph $G_{a2}$, thereby allowing the two states of the QL bit to have a frequency gap.
@@
Figure 4: Ensemble-averaged Kuramoto order parameter as a function of time predicted for the system of oscillators and compared to the purity of the emergent QL state with largest eigenvalue. When the coupling, $K$, is sufficiently large, the system equilibrates to limit cycle dynamics, indicating a synchronized classical system. The state space concomitantly evolves to an equilibrium of purity-preserving dynamics. (a) $K=100$, averaged over 100 iterations. (a) $K=200$, averaged over 100 iterations. (a) $K=400$, averaged over 500 iterations.
@@
In the section, we have generalized our prior model, allowing any phase for the edges. This enabled us to consider explicitly the extent of phase synchronization in the classical network, which was shown to control the mixedness of the emergent QL states. Hence, this property of the variables defined by the classical network influences an important property of the states—phase coherence of the basis states is essential for sustaining stable superposition states. Depending on the properties of the of the classical phase oscillator network, the QL states can exhibit purity-preserving dynamics, or they can be completely mixed and undergo classical dynamics, or any case in between. When the oscillator network is synchronized, that is, the phases are coherent, then the QL states have high purity.
V.5 Decoherence
We have focused on synchronization of the classical network to produce robust QL states. We now consider a classical network where the couplings are weak compared to the standard deviation of frequency offsets. We thereby ensure that the system cannot synchronize (see Sec. V.B). We set the state preparation so that all the oscillators are initially synchronized. We hypothesize that, with these model parameters, a QL state that initially has high purity, will become more mixed with time as the network de-synchronizes. That is, the opposite dynamics to those illustrated in Fig. 4 will ensue.
An obvious example is a network of uncoupled oscillators with a frequency distribution. Starting in-phase, free evolution of the oscillators will dephase the order parameter and states in the network. The mechanism for such decoherence is similar to that for decoherence of a qubit[22]. Each coefficient in the chosen measurement basis of the effective state acquires a phase offset from the mean as a function of time. Let’s write this as $c_{m}^{i}=e^{i\zeta_{m}t}$. Then an off-diagonal density matrix element in this basis evolves as
$\rho^{i}_{mn}(t)=\sum_{\text{prep}}e^{i(\zeta_{n}-\zeta_{m})t}.$ (35)
In the average, the phases in the emergent states evolve randomly because of the complexity of the underlying classical system, which leads to loss of phase coherence in these off-diagonal density matrix elements (they decay). This decoherence is independent of the measurement basis, unless we set the system up so that there is a preferred basis.
@@ V.5 Decoherence
Figure 5: Results of calculations of an oscillator system comprising the Cartesian product of two QL bits, using the same procedures and parameters as the previous calculations, except here we set the coupling to be weak, $K=30$ and the initial condition sets all the oscillators in phase. The averaging is carried out over a single graph. (a) The order parameter, initially 1, indicating the initial synchronization condition, decays as the oscillators become out of phase. (b) As the oscillators dephase, the purity of the emergent state decays to an equilibrium mixed state. (c) Concomitantly, off-diagonal elements of the density matrix in the product basis decay to zero, highlighting loss of superpositions.
@@
In Fig. 5 we show results of calculations of an oscillator system comprising the Cartesian product of two QL bits, using most of the same procedures and parameters as the previous calculations. Two key parameters were changed: the coupling was set to be weak, $K=30$, and the initial condition sets all the oscillators in phase. The average is over 500 realizations of state preparation.
@@
V.6 QL systems versus quantum systems
Here we speculate on what these studies of QL systems might inspire in terms of probing open questions about quantum systems.
Hidden variable theories in quantum mechanics aim to enable a physical interpretation of quantum mechanics, in particular to address questions such as how measurement works, what is the relationship between quantum and classical theories, and so on[54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. The main scope of hidden variable theories is to ask whether there is a deterministic system with predetermined outcomes prescribed for its properties, but this picture is hidden to us by the ways we measure the state vector. The motivation is that the hidden picture would clarify some of the mysteries that underpin the quantum theory. The no-hidden-variables theorems (e.g. Bell’s theorem) show how this is not the case[55].
The QL model gives a framework for examining hidden variables models in a new way. The reader will immediately object because Bell’s theorem, the Kochen-Specker theorem, and so on, are theorems that prove hidden variables cannot explain the perplexing outcomes of measurements on quantum systems. But, this result is consistent with how we find the interplay of the classical system and the QL states in the present work. It appears that correlations exhibited by the QL states are also built into the classical network. While the connections are preliminary at this stage, it does seem to be a topic worthy of additional scrutiny.
The main point is that as long as the classical system is constructed according to the same rules governing the construction of the quantum states—which is ensured by the QL graph—then the classical system and quantum states are bound such that we cannot arbitrarily assert maps on the QL states. That is, even if hidden variables do underlie a quantum system, they do not allow us to circumvent or rationalize observations that depend on quantum correlations. As an example, let’s consider the well-known no-cloning theorem[68].
Theorem 1
There is no unitary operator $U$ on $\mathcal{H}\otimes\mathcal{H}$ such that for all normalized states $|a\rangle_{A}$ and $|b\rangle_{B}$ in $\mathcal{H}=\mathcal{H}_{A}=\mathcal{H}_{B}$
$U\Big{(}|a\rangle_{A}|b\rangle_{B}\Big{)}=e^{i\alpha(a,b)}|a\rangle_{A}|a\rangle_{B}$
where $\alpha\in\mathbb{R}$.
Recall that if $|a\rangle_{A}$ is a state of the QL graph $G_{A}$ and $|b\rangle_{B}$ is a state of the QL graph $G_{B}$, then $|a\rangle_{A}\otimes|b\rangle_{B}$ is a state of the Cartesian product of those graphs, $G_{A}\Box G_{B}$. The adjacency matrix of $G_{A}\Box G_{B}$ can be drawn in block form in such a way that the inherited edge structure from $G_{B}$ is highlighted on the diagonal:
$A_{A\Box B}=\begin{bmatrix}A_{B}&*&*&*\\ *&A_{B}&*&*\\ *&*&A_{B}&*\\ *&*&*&A_{B}\end{bmatrix}$ (36)
and all the other blocks, written as []delimited-[][*][ ∗ ] contain edge structure inherited from $G_{A}$.
When translated to graphs of QL states, the no-cloning theorem says that there is no unitary matrix that can transform $A_{A\Box B}$ such that, to an overall phase factor, diagonal blocks $[A_{B}]$ are mapped to copies of $[A_{A}]$, while leaving all the blocks []delimited-[][*][ ∗ ] unchanged.
This can be shown to be the case generally, but it is somewhat tedious. In any case, here we need to consider the unitary transformation that we can implement on our network of phase oscillators, which is given by $\Phi$, Eq. 13. The transformation is to make a change to the phase of each oscillator in the network. We just need to consider the effect of the map on a few elements of the matrix to understand why cloning the state $|a\rangle_{A}$ is impossible. The map $\Phi^{-1}A_{A\Box B}\Phi$ introduces a phase rotation $e^{i(\theta_{2}-\theta_{1})}$ in the first off-diagonal element of the top-left $[A_{B}]$ block, and a rotation $e^{i(\theta_{n}-\theta_{m})}$, where $m=n_{0}^{2}+1$ and $n=n_{0}^{2}+2$, in the analogous element of the second $[A_{B}]$ block. These blocks must be changed identically by the map, so $\theta_{2}-\theta_{1}=\theta_{n}-\theta_{m}$, or equivalently $\theta_{m}-\theta_{1}=\theta_{n}-\theta_{2}$.
Now let’s consider the first []delimited-[][*][ ∗ ] block, which we require to be unchanged by the map because $|a\rangle_{A}$ should not be changed during the cloning except by a constant phase. The first two elements of the top row of []delimited-[][*][ ∗ ] are $e^{i(\theta_{m}-\theta_{1})}$ and $e^{i(\theta_{n}-\theta_{1})}$. Thus we require that $\theta_{m}-\theta_{1}=\theta_{n}-\theta_{1}$, that is, $\theta_{m}=\theta_{n}$. Clearly this us impossible to achieve if we want the $[A_{B}]$ blocks to be transformed. Therefore, even though the underlying network encoded by the graph is classical, there is no way of changing the phases of its constituent phase oscillators that allows a QL bit state to be cloned.
Another program in quantum theory is to understand how superpositions might collapse under dynamics governed by a modified version of the Schrödinger equation[69, 70, 71, 3]. Here the challenge is to understand how superpositions are lost, which is impossible under the linear evolution of the Schrödinger equation. Solving that problem would explain why macroscopic systems are not found in superposition and the theory would thus demystify the quantum to classical transition. One proposal is that the Schrödinger equation is modified so that at the micro-scale it appears unchanged from our usual formulation, but at longer length scales a nonlinear, stochastic feature emerges[70]. This new feature acts to collapse superpositions. Details of the nonlinear feature are hidden when we measure the state vectors.
The physical mechanisms for dynamical evolution of QL states suggests a connection might be made to this topic, in particular, it might inspire hypotheses to explain a mechanism for how superposition states can spontaneously collapse[3, 71, 70, 72]. Commonly, researchers incorporate a term into the Schrödinger equation that has no effect for typical quantum states, but for distant superpositions a stochastic, nonlinear term emerges that can initiate collapse of superpositions. The QL model considered in the present work appears to produce a similar result. We posited a somewhat different mechanism than proposed in prior important papers[3, 71, 70, 72] because in the present work the nonlinearity comes from the dynamics of the explicit classical system underlying the state space and it only controls phases of the underlying variables. The crucial parameter that impacts the states is the phase coherence of the classical network. However, the net result would be a modification of the Schrödinger equation prediction of dynamics in the state space. Indeed, we have noted how, when the oscillator network is poorly synchronized, it acts as a kind of bath for the state space. Thereby, it introduces a mechanism for decoherence similar to models for open quantum systems[73], where a system interacts with an external bath comprising many degrees of freedom.
V.7 Discussion
@@
The model studied here suggests a way to understand a smooth transition in the state space from QL entangled states to mixed states. This would be observed if the propensity of the classical network to synchronize diminishes as the size of the system increases. An open question, therefore, is whether (and how) there can be a distance dependence governing the decoherence so that distant superpositions naturally collapse. For example, we can imagine that the noise that gives the oscillators the frequency distribution is correlated on molecular length scales, but uncorrelated on longer length scales. This setup might serve to collapse distant superpositions, but preserve superpositions among closely-spaced systems.
@@ VI Conclusions
We studied a specific example of a large classical system that maps, via a graph, to a state space that we call quantum like (QL) because it mimics the kinds of states typical of quantum systems. The classical system is a network of coupled phase oscillators, of the kind extensively studied in a wide variety of contexts. Such networks of coupled phase oscillators evolve according to the nonlinear dynamics associated with synchronization of the oscillator phases. We reported numerical simulations for dynamics of this system, starting from a state preparation produces a random set of initial oscillator phases. The network topology was blueprinted by a graph that generates the QL states. We used the phases of the classical system to weight the edges of the QL state graph at various time points during the dynamics, then mapped those graphs to the QL state space. Studying the state space, we constructed the density matrix in an effective basis of two coupled QL bits and showed that the mixedness of the emergent state diminished as the underlying classical system synchronized.
@@
Key conclusions of the work are that nonlinear dynamics in the classical system controls the dynamics in the state space. The dynamics can help quantum-like states be robust to noise and disorder, or they can act like an intrinsic environment to produce decoherence in the state space. We showed numerical studies of both scenarios. Nevertheless, these dynamics convey linear maps in the state space (Prop. 1). This setup suggests a mechanism by which underlying variables act to collapse entangled states in the QL system. It is speculated that the concept might inspire a plausible mechanism that, with suitable development, might explain the quantum-to-classical transition. Further, we compared QL states to quantum states. We found that during decoherence of the QL states, the off-diagonal density matrix elements decay and that this decay can be observed in any basis we choose for measurement. More surprisingly, we showed that a no-cloning theorem (that is, a state of a QL bit cannot be copied) applies not only to the QL states, but also to the underlying classical system.